# Attitude Trajectories Stabilizable by Linear Control

### v1

@@ III Linear constant difference attitude control
Dynamical equations (17) and (19), together with the constant angle constraint $\dot{e}_{0}=0$, form the starting point for the following stability analysis. The governing equations for $\vec{e}_{v}$ and $\vec{w}$ may be written in state-space form as
$\begin{bmatrix}\dot{\vec{e}}_{v}\\ \dot{\vec{w}}\end{bmatrix}$ $=\begin{bmatrix}(e_{0}-1)\vec{w}+\vec{e}_{v}\vec{e}_{v}^{T}\vec{w}+\vec{e}_{v}\times\vec{w}\\ |\vec{w}|^{2}\vec{e}_{v}/(1+e_{0})\end{bmatrix}.$ (20)
@@
We may linearize the above dynamics, and use the constant difference constraint $\vec{e}_{v}^{T}\vec{w}=0$, to obtain
$\begin{bmatrix}\dot{\tilde{e}}_{v}\\ \dot{\tilde{w}}\end{bmatrix}$ $=\begin{bmatrix}(e_{0}-1)\tilde{w}+\vec{e}_{v}\vec{w}^{T}\tilde{e}_{v}+\vec{e}_{v}\vec{e}_{v}^{T}\tilde{w}+\tilde{e}_{v}\times\vec{w}+\vec{e}_{v}\times\tilde{w}\\ (|\vec{w}|^{2}\tilde{e}_{v}+2\vec{e}_{v}\vec{w}^{T}\tilde{w})/(1+e_{0})\end{bmatrix}$
$=A\begin{bmatrix}\tilde{e}_{v}\\ \tilde{w}\end{bmatrix},$ (21)
@@
where $\tilde{e}_{v}$ and $\tilde{w}$ are perturbations in the vector part $\vec{e}_{v}$ of the attitude error, and the (half) angular velocity $\vec{w}$, and
$A$ $=\begin{bmatrix}\vec{e}_{v}\vec{w}^{T}-[\vec{w}\times]&(e_{0}-1)1_{3\times 3}+[\vec{e}_{v}\times]+\vec{e}_{v}\vec{e}_{v}^{T}\\ \dfrac{|\vec{w}|^{2}1_{3\times 3}}{1+e_{0}}&\dfrac{2\vec{e}_{v}\vec{w}^{T}}{1+e_{0}}\end{bmatrix}.$ (22)
@@
The trace of $A$ above may be immediately verified to be equal to zero after using (14). This implies that all the eigenvalues are on the imaginary axis, or at least one of them has a positive real part and the system is unstable.
@@
Although the $6\times 6$ matrix $A$ in (22) looks quite formidable, it may be simplified significantly by applying the following orthogonal transformation. Let
$C$ $=\begin{bmatrix}\dfrac{\vec{e}_{v}}{|\vec{e}_{v}|}&\dfrac{\vec{w}\times\vec{e}_{v}}{|\vec{e}_{v}||\vec{w}|}&\dfrac{\vec{w}}{|\vec{w}|}\end{bmatrix};$ (23)
then,
$A^{\prime}=\begin{bmatrix}C^{T}\\ &C^{T}\end{bmatrix}A\begin{bmatrix}C\\ &C\end{bmatrix}$ (24)
@@
where, $c=e_{0}$ and $s=|\vec{e}_{v}|$, hinting trigonometric ratios cosine and sine for the scalar and vector parts of the error quaternion $\check{e}$. For instance, $c$ would be the cosine of half the error angle. Note that empty entries in the matrix $A$ indicate a zero.
Since $A$ and $A^{\prime}$ are related through the orthogonal transformation in (24), they share the same characteristic polynomial. The characteristic polynomial $\det(\lambda 1_{6\times 6}-A^{\prime})$ for the matrix $A^{\prime}$ in (27) may be evaluated to be equal to
$f_{A}(\lambda)=f_{A^{\prime}}(\lambda)=\frac{\lambda^{6}}{|\vec{w}|^{6}}+\frac{(3-2e_{0}+e_{0}^{2})}{(1+e_{0})}\frac{\lambda^{4}}{|\vec{w}|^{4}}$
$\quad+\frac{(1-e_{0})(4+e_{0}+3e_{0}^{2})}{(1+e_{0})^{2}}\frac{\lambda^{2}}{|\vec{w}|^{2}}+\frac{2e_{0}^{2}(1-e_{0})^{2}}{(1+e_{0})^{3}}.$ (28)
An immediate observation is that the characteristic polynomial $f_{A}(\lambda)$ contains only even powers of $\lambda$ so roots are point symmetric about the origin. Multiplying $f_{A}$ with $(1+e_{0})^{3}$ and substituting $\lambda^{\prime}$ for $(1+e_{0})\lambda^{2}/|\vec{w}|^{2}$, $f_{A}(\lambda)$ may be simplified to the following cubic in terms of $\lambda^{\prime}$
$\lambda^{\prime 3}+(3-2e_{0}+e_{0}^{2})\lambda^{\prime 2}+(1-e_{0})(4+e_{0}+3e_{0}^{2})\lambda^{\prime}$
$\quad+2e_{0}^{2}(1-e_{0})^{2}.$ (29)
Since the roots of a cubic polynomial can be obtained in closed form, it is therefore possible to analytically compute the eigenvalues of the sixth order system in (21). This rather remarkable outcome is a consequence of the special structure associated with the dynamical equations (17) and (19) of the system under consideration (21), and their linearization about the constraint $\vec{e}_{v}^{T}\vec{w}=0$ in (14). In particular, the elements of $A$ may all be expressed in terms of standard Euclidean vector operations such as the scalar and vector product, and the projection operator. And the characteristics of systems described using such Euclidean vector expressions are invariant upon the use of orthogonal transformations. The specific orthogonal transformation $C$ used in (24) aligns the $x$-axis along $\vec{e}_{v}$ and the $z$-axis along $\vec{w}$.
@@
$\!\!\!\!\Delta$ $=e_{0}^{6}+16e_{0}^{5}+70e_{0}^{4}+56e_{0}^{3}-151e_{0}^{2}+168e_{0}-112$ (30)
is positive (the trivial factors consist of twice repeated roots at $-1$ and $+1$). Within the range [1, 1]11[-1,\;1][ - 1 , 1 ] for $e_{0}$, the discriminant is positive when $e_{0}\gtrsim 0.85$ (error angle less than $\approx 31.7^{\circ}$) and negative otherwise. Furthermore, the coefficients in $f_{A}$ are nonnegative for all possible values of $e_{0}$, and the product of the coefficients of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient in (29), allowing us to conclude that the roots of (29) are always in the left half of the complex plane (that is, their real parts are negative). When they are purely real negative, that leads to purely imaginary roots in (28), and the system is marginally stable despite the lack of active negative feedback induced stabilization. When the roots of (29) are complex for $e_{0}\lesssim 0.84$ (error angle greater than $\approx 32.9^{\circ}$), that leads to (28) having roots with both positive and negative real parts and consequent instability.

### v2

@@ III-B Linearized (Perturbation) Dynamics
Dynamical equations (19) and (21), together with the constant angle constraint $\dot{e}_{0}=0$, define the nominal motion for $\check{e}$ and $\vec{\omega\mkern-1.0mu}$ in the following perturbation and stability analysis. They may be written in (nonlinear) state-space form as
$\begin{bmatrix}\dot{\vec{e\mkern-1.0mu}}_{\text{v}}\\ \dot{\vec{w\mkern-1.0mu}}\end{bmatrix}$ $=\begin{bmatrix}(e_{0}-1)\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{w\mkern-1.0mu}\\ |\vec{w\mkern-1.0mu}|^{2}\vec{e\mkern-1.0mu}_{\text{v}}/(1+e_{0})\end{bmatrix}$. (22)
@@
We may linearize the above dynamics, and use the constant difference constraint $\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}=0$, to obtain
$\begin{bmatrix}\dot{\tilde{e}}_{\text{v}}\\ \dot{\tilde{w}}\end{bmatrix}$ $=A\begin{bmatrix}\tilde{e}_{\text{v}}\\ \tilde{w}\end{bmatrix}$, (23)
@@
where $\tilde{e}_{\text{v}}$ and $\tilde{w}$ are perturbations in the vector part $\vec{e\mkern-1.0mu}_{\text{v}}$ of the attitude error, and the (half) angular velocity $\vec{w\mkern-1.0mu}$, (so, the corresponding total quantities are $\vec{e\mkern-1.0mu}_{\text{v}}+\tilde{e}_{\text{v}}$ and $\vec{w\mkern-1.0mu}+\tilde{w}$) and
$A$ $=\begin{bmatrix}\vec{e\mkern-1.0mu}_{\text{v}}\vec{w\mkern-1.0mu}^{\text{T}}-[\vec{w\mkern-1.0mu}\times]-\dfrac{\vec{w\mkern-1.0mu}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}}{e_{0}}&(e_{0}-1)1_{3\times 3}+[\vec{e\mkern-1.0mu}_{\text{v}}\times]+\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\\ \dfrac{|\vec{w\mkern-1.0mu}|^{2}1_{3\times 3}}{1+e_{0}}+\dfrac{|\vec{w\mkern-1.0mu}|^{2}e_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}}{(1+e_{0})^{2}e_{0}}&\dfrac{2\vec{e\mkern-1.0mu}_{\text{v}}\vec{w\mkern-1.0mu}^{\text{T}}}{1+e_{0}}\end{bmatrix}$. (24)
@@
We have substituted $\tilde{e}_{0}$ with $\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\tilde{e}_{\text{v}}/e_{0}$ in the above expression. The trace of $A$ above may be immediately verified to be equal to zero after using (13). This implies that all the eigenvalues are on the imaginary axis, or at least one of them has a positive real part and the system is unstable.
@@
Although the $6\times 6$ matrix $A$ in (24) looks quite formidable, it may be simplified significantly by applying the following orthogonal and scaling transformations. Let
$C$ $=\begin{bmatrix}\dfrac{\vec{e\mkern-1.0mu}_{\text{v}}}{|\vec{e\mkern-1.0mu}_{\text{v}}|}&\dfrac{\vec{w\mkern-1.0mu}\times\vec{e\mkern-1.0mu}_{\text{v}}}{|\vec{e\mkern-1.0mu}_{\text{v}}||\vec{w\mkern-1.0mu}|}&\dfrac{\vec{w\mkern-1.0mu}}{|\vec{w\mkern-1.0mu}|}\end{bmatrix}$; (25)
then, it is a simple matter to verify that
$C^{\text{T}}C$ $=1_{3\times 3},$ $C^{\text{T}}\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}C$ $=\begin{bmatrix}|\vec{e\mkern-1.0mu}_{\text{v}}|^{2}&&\\ &0&\\ &&0\end{bmatrix}$, (26)
$C^{\text{T}}\vec{e\mkern-1.0mu}_{\text{v}}\vec{w\mkern-1.0mu}^{\text{T}}C$ $=\begin{bmatrix}&&|\vec{e\mkern-1.0mu}_{\text{v}}||\vec{w\mkern-1.0mu}|\\ &0&\\ 0&&\end{bmatrix}$, $C^{\text{T}}\vec{w\mkern-1.0mu}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}C$ $=\begin{bmatrix}&&0\\ &0&\\ |\vec{e\mkern-1.0mu}_{\text{v}}||\vec{w\mkern-1.0mu}|&&\end{bmatrix}$,
$C^{\text{T}}[\vec{e\mkern-1.0mu}_{\text{v}}\times]C$ $=\begin{bmatrix}0\\ &&-|\vec{e\mkern-1.0mu}_{\text{v}}|\\ &|\vec{e\mkern-1.0mu}_{\text{v}}|&\end{bmatrix}$, $C^{\text{T}}[\vec{w\mkern-1.0mu}\times]C$ $=\begin{bmatrix}&-|\vec{w\mkern-1.0mu}|\\ |\vec{w\mkern-1.0mu}|&&\\ &&0\end{bmatrix}$.
We now rotate the state matrix $A$ in (24) through $C$
$A^{\prime}=\begin{bmatrix}C^{\text{T}}\\ &C^{\text{T}}\end{bmatrix}A\begin{bmatrix}C\\ &C\end{bmatrix}$ (27)
@@
where, $c=e_{0}$ and $s=|\vec{e\mkern-1.0mu}_{\text{v}}|$ are the cosine and sine of the error half angle, and $w=|\vec{w\mkern-1.0mu}|$. Note that empty entries in a matrix indicate a zero.
Since $A$ and $A^{\prime}$ are related through the orthogonal transformation in (27), they share the same characteristic polynomial. The characteristic polynomial $\det(\lambda 1_{6\times 6}-A^{\prime})$ for the matrix $A^{\prime}$ in (30) may be evaluated to be equal to (see the appendix -A for a derivation)
$f_{A}(\lambda)=f_{A^{\prime}}(\lambda)=\frac{\lambda^{6}}{|\vec{w\mkern-1.0mu}|^{6}}+\frac{(1+3e_{0}-e_{0}^{2}-e_{0}^{3})}{e_{0}(1+e_{0})}\frac{\lambda^{4}}{|\vec{w\mkern-1.0mu}|^{4}}$
$\quad+\frac{(1-e_{0})(3+3e_{0}-2e_{0}^{2})}{(1+e_{0})^{2}}\frac{\lambda^{2}}{|\vec{w\mkern-1.0mu}|^{2}}+\frac{2e_{0}(1-e_{0})^{2}}{(1+e_{0})^{3}}.$ (31)
An immediate observation is that the characteristic polynomial $f_{A}(\lambda)$ contains only even powers of $\lambda$ so roots are point symmetric about the origin. Multiplying $f_{A}$ with $(1+e_{0})^{3}$ and substituting $\lambda^{\prime}$ for $(1+e_{0})\lambda^{2}/|\vec{w\mkern-1.0mu}|^{2}$, $f_{A}(\lambda)$ may be simplified to the following cubic in terms of $\lambda^{\prime}$
$\lambda^{\prime 3}+(1/e_{0}+3-e_{0}-e_{0}^{2})\lambda^{\prime 2}$
$\quad+(1-e_{0})(3+3e_{0}-2e_{0}^{2})\lambda^{\prime}+2e_{0}(1-e_{0})^{2}.$ (32)
Since the roots of a cubic polynomial can be obtained in closed form, it is therefore possible to analytically compute the eigenvalues of the sixth order system in (23). This rather remarkable outcome is a consequence of the special structure associated with the dynamical equations (19) and (21) of the system under consideration (23), and their linearization about the constraint $\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}=0$ in (13). In particular, the elements of $A$ may all be expressed in terms of standard Euclidean vector operations such as the scalar and vector product, and the projection operator. And the characteristics of systems described using such Euclidean vector expressions are invariant upon the use of orthogonal transformations. The specific orthogonal transformation $C$ used in (27) aligns the $x$-axis along $\vec{e\mkern-1.0mu}_{\text{v}}$ and the $z$-axis along $\vec{w\mkern-1.0mu}$.
@@
$\Delta$ $=1+5e_{0}-8e_{0}^{2}+4e_{0}^{3}$ (33)
is positive (the trivial factors consist of a root at 1-1 and eight at +1+1). Within the range [1, 1][-1,\;1] for $e_{0}$, the discriminant is positive when $e_{0}\gtrsim-0.16$ (error angle less than $\approx 99^{\circ}$) and negative otherwise. Furthermore, the coefficients in $f_{A}$ are nonnegative for all possible values of $e_{0}$, and the product of the coefficients of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient in (32), allowing us to conclude that the roots of (32) are always in the left half of the complex plane (that is, their real parts are negative). When they are purely real negative, that leads to purely imaginary roots in (31), and the system is marginally stable despite the lack of active negative feedback induced stabilization. When the roots of (32) are complex for $e_{0}\lesssim-0.16$ (error angle $\gtrsim 99^{\circ}$), that leads to (31) having roots with both positive and negative real parts and consequent instability.